OpenStudy (anonymous):

Subtract and simplify: x + y over 6 minus quantity 2x + y over 10

OpenStudy (anonymous):

\[(x+y)/6 - (2x+y)/10\]

OpenStudy (anonymous):

\[(10x +10y)/60 - (12x+6y)/60\]

sam (.sam.):

\[\frac{x+y}{6}-\frac{2x+y}{10}\] Make denominator same, \[\frac{(x+y)\times 10}{6\times 10}-\frac{(2x+y)\times 6}{10\times 6}\] \[\frac{10x+10y}{60}-\frac{(12x+6y)}{60} \] Now that the denominator is equal you can combine them, \[\frac{(10x+10y)-(12x+6y)}{60}\] Can you simplify it now?

OpenStudy (anonymous):